OpenStudy (anonymous):

Induction Question Wording? I copied the question below. The problem is easy to solve but the wording make its a bit confusing. It says 30 degree angle to the plane. This makes me think that the angle between the normal vector of the plane and the magnetic field is 60 degrees, not 30 degrees. However, the solution shows that 30 degrees is used in Faraday's law to get the induced emf. Can someone please explain to me why this is what the solution shows or do you think the solution is incorrect? A wire loop with an area of 0.020 m2 is in a magnetic field of 0.30 T directed at 30° angle to the plane of the loop. If the field drops to zero in 45 ms, what is the average induced emf in the loop?

OpenStudy (anonymous):

Ok so here is the solution you are right since the plane is making an angle of 30 degrees, the normal makes an angle of 60 degrees and hence the net flux throught the loop = .3 X .02 X cos (60) = 3 X 10^(-3) Tm^2 hence the average emf induced = \[e = \frac{\Delta \phi }{\Delta t} = \frac{3 \times 10^{-3}}{45 \times 10^{-3}} = \frac{1}{15}V\]

OpenStudy (anonymous):

Oh alright. Then maybe the solution for the problem is incorrect. Thank you!